Malome is presently half his brother's age. Five years ago his age and his brother's age added up to 26. What is Malome's present age.?
Found 2 solutions by stanbon, ewatrrr:
Answer by stanbon(75887)

Malome is presently half his brother's age. Five years ago his age and his brother's age added up to 26. What is Malome's present age.?
--------------
Equations:
m = (1/2)b
(m-5)+(b-5) = 26
----------------------
Substitute for "m" and solve for "b":
(1/2)b -5 + b-5 = 26
(3/2)b - 10 = 26
(3/2)b = 36
b = (2/3)36
b = 24 yrs (Brother's age now)
----------------
Solve for "m"
m = (1/2)b
m = 12 yrs (Malcolm's age now)
================================
Cheers,
Stan H.
=============

Answer by ewatrrr(24785)

 
Hi
Malome is presently half his brother's age
Let x and 2x represent Malome and his brother's ages NOW
Question states***5 years ago...
(x-5) + (2x - 5) = 26
3x = 36
x = 12, Malome's age NOW. His brother is 24. (2*12)
CHECKING our Answer***5 years ago...
7 + 19 = 26